OpenStudy (anonymous):

The Wronskian of 2 functions f and g is (x^2)exp(x) . Find g(x) if f(x) = x Note : X square exponential x

OpenStudy (anonymous):

hmm i thought wronskian was something else but there is no f or g in \[x^2e^x\]

OpenStudy (jamesj):

By definition the Wronskian W of f and g is the determinant of this matrix f' g' f g I.e., W(f,g) = f'g - fg'. Use the information in your problem to write down a differential equation for g

OpenStudy (anonymous):

i will be quiet and learn, but i thought it was \[gf'=fg'\]

OpenStudy (anonymous):

ok now i feel like a moron, i misinterpreted the question completely

OpenStudy (anonymous):

JamesJ - But after that you will get a non homogeneous first order diff equation . Am i right? I'm stuck there Satelite - Thanks anyway

OpenStudy (lalaly):

what was the equation u got?

OpenStudy (jamesj):

W(f,g) = f'g - fg'. Thus x^2e^x = g - xg' or in more standard form, g' - (1/x)g = -xe^x Now you can solve that with an integrating factor.

OpenStudy (anonymous):

xg'+g=(x^2)(exp(x)) x square exponential x

OpenStudy (lalaly):

notice the left side is the product rule (xg)' = x^2e^x integrate both sides

OpenStudy (anonymous):

Can someone help me out with the integrating factor formula . Is it the same formula we used for non homogenoeous?

OpenStudy (anonymous):

x . g(x) = integral of <x^2 . e^x Am i right?

OpenStudy (jamesj):

@lalady, not quite, because of the minus sign. Or at least not in the correct equation. @ghan: watch this lecture: http://ocw.mit.edu/courses/mathematics/18-03-differential-equations-spring-2010/video-lectures/lecture-3-solving-first-order-linear-odes/ Using Integrating Factors is a very important basic tool for ODEs and worth learning. The lecture does it well.

OpenStudy (lalaly):

oh in his equation there was no minus, im sorry

OpenStudy (anonymous):

I made a mistake . Sorry guys